the product of two whole numbers is 144. If one number is four times the other number, what are the numbers?
Answer by t0hierry(194)

Call one number a and the other b. b = 4a. Thus the product of the two is 4a^2
Now write 4a^2 = 144
This gives you
a^2 = 36
or a = 6
and b = 24
Check this is right by computing ab = 6*24 = 144